The Federal Reserve Bank purchases $X worth of securities and as a result, total checkable deposits in the economy increases by $3,500? Using a required reserve ratio of 6% and the simple multiplier formula, find X.

Group of answer choices

$175

$190

$210

$245
Answer #1

Answer

The correct answer is (c) $210

Formula :

Simple money multiplier(m) = 1/rr where rr = required reserve ratio = 6% = 0.06 => m = 1/0.06

Total Checkable Deposit(D) = m*R

D = (1/0.06)R

AR => AD= 0,06

where R = Increase in reserves and here Federal Reserve Bank purchases of X(Open market purchase) will increase reserves(R) by X => AR= X. Also Total checkable deposit(D) increases by 3500 => AD = 3500

Hence using above information and formula we have :

AR AD = 0,06

=> 3500 = X/0.06

=> X = 0.06*3500 = 210

Thus, X = 210

Hence, the correct answer is (c) $210